OpenStudy (anonymous):

perimeter of 18ft by area of 8ft

OpenStudy (anonymous):

A rectangle?!

OpenStudy (anonymous):

what is the area of a rectangle?

OpenStudy (anonymous):

sorry i meant is it a rectangle.

OpenStudy (anonymous):

square

OpenStudy (anonymous):

it cannot be a square.

OpenStudy (anonymous):

the question is"we need a rock garden with a perimeter of 18ft and the area of 8ft

OpenStudy (radar):

Let l=length Let w=width lw=8 sq ft (area given) 2l+2w=18 ft (perimeter given) l=8/w from the first equation above. Now subsitute in second equation for l 2(8/w)+2w=18 16/w+2w=18 16+2w^2=18w multiplying all terms by w \[2w ^{2}-18w+16=0\] \[w ^{2}-9w+8=0\] Dividing all terms by 2 Now factor (w-8)(w-1)=0 w=8 or w=1 when w is 8, l is 1 when w is 1, l is 8 your rectangle is 8 by 1

OpenStudy (radar):

Did you follow redboi06?